solve:- (2x 2 -9x+4) 1/2 + 3(2x-1) 1/2 =(2x 2 +21x-11) 1/2

solve:-


(2x2-9x+4)1/2 + 3(2x-1)1/2 =(2x2+21x-11)1/2

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sayam

(2x2-9x+4)1/2 + 3(2x-1)1/2 =(2x2+21x-11)1/2

square

(2x2-9x+4) + 9(2x-1)  +6(2x2-9x+4)1/2(2x-1)1/2 =(2x2+21x-11)

simplify

 (2x2-9x+4)1/2 =(2x-1)1/2

square again

(2x2-9x+4) =(2x-1) 

 (2x2-11x+5) =0

now solve

 x= (11 ± √41) /4


 

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin